# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v3

@@ 2 NDGM and convergence analysis
A common limitation in the current and subsequent sections is that we do not analyze the dynamics of the neural network training process as it is often conducted in practice. This is necessitated by the fact that neural networks often involve non-differentiable layers like dropout and max pooling, which are beyond the scope of our study. We do not reiterate this limitation in every subsequent section.
@@
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks & Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.
@@ 3 NDGM and the LASSO penalty
Zero weights are crucial for model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and inducing sparsity by penalizing the $L_{1}$ norm of model weights is a common approach. While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017). This expectation is not always met, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli. Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li & Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
@@
We show that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model. In particular, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty. We first analytically demonstrate the unreliability for the simplest LASSO problem,
@@
$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
@@
See Appendix B for details ∎
Equation (8) suggests that with the same learning rate and initialization, a larger $\lambda_{1}$ can lead to a NDGM solution with a larger $L_{1}$ norm, defeating the purpose of the LASSO penalty. To illustrate this, we note that in the 2D case, with weights initialized as $\beta_{0}=[0.5053,0.5053]$, $\alpha=0.01$, and $\lambda_{1}=1$, the NDGM sequence oscillates between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$. However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha_{1}$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Therefore at convergence, the $L_{1}$ norm is at most 0.0106 for $\lambda_{1}=1$, but at least 0.9894 for $\lambda_{1}=100$. Furthermore, reducing the step-size to $\alpha=0.001$ after apparent convergence does not resolve this paradox. With $\lambda_{1}=1$, the sequence oscillates between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$, while with $\lambda_{1}=100$, it oscillates between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$. Appendix C shows that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
@@
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from (Boyd et al., 2003) (Section 2 on the Constant Step Size) that
@@
Conventional wisdom in deep learning theory suggests that different NDGM variants, such as NDGM with momentum and RMSProp, converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of (Goodfellow et al., 2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made). However, we prove in Appendices D and E that these claims are incorrect even for the toy LASSO problem described in (7). Furthermore, our analysis reveals two interesting facts. First, that the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis. Second, there is a phase transition between the RMSProp and vanilla NDGM training curves characterized by a change in the relative behavior of the training curves of the two algorithms as $\lambda_{1}$ is increased from a value which is much smaller than 1 to a value much larger than 1, which cannot be explained without accounting for the non-differentiability of the $L_{1}$ regularizer. These findings highlight the importance of incorporating the non-differentiability of regularizers like the $L_{1}$ norm when analyzing the behavior of NDGM variants. Failing to account for non-differentiability results in an incomplete understanding of the optimization dynamics. Consequently, it remains unclear how the convergence analysis of algorithms like Adam and RMSProp for differentiable loss functions, as often performed in studies such as (Reddi et al., 2019; Défossez et al., 2020; Zhang et al., 2022), relates to the dynamics of these algorithms’ convergence in non-differentiable settings involving regularizers like the $L_{1}$ norm or activation functions like the ReLU.
@@
In their seminal work on network pruning, (Han et al., 2015b) propose using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s. To challenge this suggestion’s general applicability, we conduct two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs. We use $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam. At the end of our runs, we make two key observations (Figure 5): 1) The $L_{1}$ norms at convergence differ vastly across the three variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence. These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may not be as straightforward as previously thought. Indeed, our results show that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 1): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do.
@@
The observation that higher $\lambda_{1}$ values do not yield more prunable parameters indicates that while soft thresholding produces sparse solutions for NDGMs, it does not find optimal solutions. This issue extends beyond NDGM to general neural network problems, where any soft thresholding solution yields sparsity but not optimality. The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of (Chen et al., 2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously. Consequently, applying the generalized ISTA formulation relying on $L-smoothness$ to neural networks necessarily leads to suboptimal solutions.
@@
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed for $L_{1}$ regularized problems (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack this capability predates the widespread adoption of deep learning (Xiao, 2009). Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016). However, there is a fundamental disparity between these frameworks: glmnet focuses on a narrow set of problems, guaranteeing optimality, while PyTorch (Paszke et al., 2017) provides flexibility in model creation, but lacks similar guarantees. Attaining a solution at a non-differentiable point is challenging, emphasizing the importance of exploring related studies and tools rather than relying on simplistic implementations when precision is crucial.
@@ 4 NDGM and the Edge of Stability
In recent years, there has been a growing interest on the topic of “unstable convergence”. The literature on the subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run. $\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^9] We define unstable convergence to mean “non-divergence to $\infty$”. Our definition is adapted from Definition 1.1 of (Arora et al., 2022). However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks. ($\alpha>\alpha^{*}$). The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
It is conjectured that neural networks demonstrate unstable convergence because the phenomenon is inconsistent with widespread presumptions in the field of optimization (Cohen et al., 2021; Ahn et al., 2024). This conjecture is incorrect for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
@@
Note that any loss function satisfying (6) has subgradients bounded above by $L$. Additionally, as outlined in Section 1 of Boyd et al. (2003), the deductions derived from the subgradient method hold equivalently for differentiable functions. This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient. Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al. (2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result. ∎
Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function. As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$. This is to be expected from Proposition 4.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
The toy LASSO problem studied in (7), for which we analytically demonstrated unstable convergence in the previous section, is an example of a non-differentiable convex loss function satisfying Proposition 4.1. An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015). The Huber loss function is non-smooth because it is once but not twice differentiable. For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (10)
@@
where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$. Figure 3 shows that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations. Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic. To emphasize this point, we provide an example of such an extrapolation in the literature which breaks down in the non-smooth case.
@@
In section 3.1 of (Ma et al., 2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near the local minimum, concluding that subquadratic growth explains the edge of stability phenomenon. However, their analysis is flawed because ReLU networks have generically non-differentiable local minima (Laurent & Brecht, 2018), where their approximation breaks down.
The Huber loss function described in equation (10) challenges their statement’s broad applicability. This overparameterized convex function has a global minimum of 0, exhibiting quadratic growth nearby. (Ma et al., 2022)’s result suggests stable convergence, but Proposition 4 guarantees unstable convergence (Figure 3). The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 in (Ma et al., 2022). In contrast, $L-smooth$ functions, which are twice differentiable by definition, exhibit subquadratic growth near their local minimum during unstable convergence. This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
@@
We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur. Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness. Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development. Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.

### v4

@@ 3 NDGM and convergence analysis
A common limitation in the current and subsequent sections is that we do not analyze the dynamics of the neural network training process as it is often conducted in practice. This is necessitated by the fact that neural networks often involve dropout layers, which leave the loss function discontinuous, let alone non-differentiable, and are thus beyond the scope of our study. We do not reiterate this limitation in every subsequent section.
Extending our results to the general neural network setting:
Our analysis in 3 focuses on the specific case of non-positive response vectors, which allows for a convex formulation. Extending these results to general single-layer neural networks with arbitrary response vectors, let alone multi-layer networks, is challenging due to the resulting non-convex and non-differentiable loss functions. Currently, there is no known way to rigorously analyze such networks in their full generality. This limitation is not unique to our work—it is a fundamental challenge in the field. Our paper’s key contribution is to rigorously demonstrate that this challenge cannot be addressed by simply extrapolating results from the analysis of smooth functions to non-smooth settings, as is often done in the literature. We specifically caution against the common practice of deriving results for $L-smooth$ networks and then claiming generality based on simulations on architectures like VGG16 or ResNet. This approach, while prevalent, fails to capture the fundamental differences between smooth and non-smooth optimization that we highlight. Our findings are crucial for the field, as they expose the limitations of current methods and underscore the need for more rigorous and appropriate techniques in analyzing ReLU networks and other non-smooth architectures.
@@
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks and Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.
@@ 4 NDGM and the LASSO penalty
Zero weights are crucial for model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and inducing sparsity by penalizing the $L_{1}$ norm of model weights is a common approach. While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder and Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou and Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017). However, this assumption is not satisfied in practice, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli. Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li and Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
@@
We demonstrate that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model. Specifically, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty. We first analytically demonstrate the unreliability for the simplest LASSO problem,
@@
$\beta_{N+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
@@
See Appendix B.1 for details ∎
Equation (8) suggests that with the same learning rate and initialization, a larger $\lambda_{1}$ can lead to a NDGM solution with a larger $L_{1}$ norm, defeating the purpose of the LASSO penalty. To illustrate this, we note that in the 2D case, with weights initialized as $\beta_{0}=[0.5053,0.5053]$, $\alpha=0.01$, and $\lambda_{1}=1$, the NDGM sequence oscillates between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$. However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Therefore at convergence, the $L_{1}$ norm is at most 0.0106 for $\lambda_{1}=1$, but at least 0.9894 for $\lambda_{1}=100$. Furthermore, reducing the step-size to $\alpha=0.001$ after apparent convergence does not resolve this paradox. With $\lambda_{1}=1$, the sequence oscillates between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$, while with $\lambda_{1}=100$, it oscillates between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$. Appendix B.2 shows via simulation that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
@@
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from Boyd et al. (2003) (Section 2 on the Constant Step Size) that
@@
Conventional wisdom in deep learning theory suggests that different NDGM variants converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of Goodfellow et al. (2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made). However, we analytically show that these claims are incorrect even for the toy LASSO problem described in (7) (see Appendices B.3 and B.4 for details). Furthermore, our analysis reveals two interesting facts which cannot be explained by the analysis of differentiable functions: 1) the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis, and 2) there is a phase transition between the RMSProp and NDGM training curves based on the magnitude of $\lambda_{1}$. These findings highlight the importance of incorporating the non-differentiability of regularizers like the $L_{1}$ norm when analyzing the behavior of NDGM variants.
@@
The seminal work on network pruning by Han et al. (2015b) proposes using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach has been widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/y7o7hw. To examine the general applicability of this method, we conducted two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs. We used $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam.
Our experiments yielded two key observations (Figure 3): 1) The $L_{1}$ norms at convergence differ substantially across the three optimizer variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence. These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may be more complex than previously understood. Indeed, our results indicate that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 2): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do. These observations raise important questions about the efficacy and generalizability of $L_{1}$ regularization for network pruning across different optimization algorithms and regularization strengths.
Refer to captionRefer to caption
Figure 3: $L_{1}$ norm of the weights as a function of iteration number for $\lambda_{1}=0.002$ (left) and $\lambda_{1}=2$ (right). In each figure, the blue, red, and green lines represent curves for RMSProp, Adam, and vanilla SGD respectively.
@@
The observation that higher $\lambda_{1}$ values do not yield more prunable parameters indicates that while soft thresholding produces sparse solutions for NDGMs, it does not find optimal solutions. This issue extends to general neural network problems, where any soft thresholding solution yields sparsity but not optimality. The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of Chen et al. (2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously. Consequently, applying the generalized ISTA formulation relying on $L-smoothness$ to neural networks necessarily leads to suboptimal solutions.
@@
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed for $L_{1}$ regularized problems (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack this capability predates the widespread adoption of deep learning (Xiao, 2009). Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer, 2016). However, there is a fundamental disparity between these frameworks: glmnet focuses on a narrow set of problems, guaranteeing optimality, while PyTorch (Paszke et al., 2017) provides flexibility in model creation, but lacks similar guarantees. Attaining a solution at a non-differentiable point is challenging, emphasizing the importance of exploring related studies and tools rather than relying on simplistic implementations when precision is crucial.
Connections to the Edge of Stability:
@@ 5 NDGM and the Edge of Stability
In recent years, there has been growing interest in the phenomenon of "unstable convergence" in neural network optimization. The literature on this subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen and Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by an intriguing observation: unlike convex quadratic forms, gradient descent on neural networks does not diverge to $\infty$ even when the learning rate exceeds $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically shown to converge "unstably", reducing non-monotonically in the long run. The value $\alpha=\alpha^{*}$ is termed the "Edge of Stability" in the literature, as it demarcates regions of "stable convergence" ($\alpha<\alpha^{*}$) from regions of "unstable convergence" ($\alpha>\alpha^{*}$).
We define unstable convergence as "non-divergence to $\infty$". This definition is adapted from Definition 1.1 and the subsequent discussion in Arora et al. (2022), which explains why using sharpness greater than $2/\eta$ is not part of the definition. We deviate from their definition due to its reliance on the Hessian existing at all points between the current and next iterate, a condition unmet in non-smooth neural networks. Importantly, the derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
It has been conjectured that neural networks exhibit unstable convergence because this phenomenon lies beyond the scope of classical optimization theory (Cohen et al., 2021; Ahn et al., 2024). However, we show that this conjecture is incorrect for a broader class of functions, specifically Lipschitz continuous, non-smooth convex functions satisfying (6), as demonstrated in the following proposition:
@@
See Appendix D.1 for details. ∎
The toy LASSO problem studied in (7) and the Huber loss function, widely used in robust regression and object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015), exemplify non-smooth convex loss functions satisfying Proposition 5.1. The LASSO problem is non-differentiable, while the Huber loss function is differentiable but non-smooth, as it is once but not twice differentiable. The Huber loss for regressing an arbitrary $50\times 1$ response vector, on a $50\times 200$ data matrix is given by
$L_{7}(\beta)$ $=\frac{1}{50}\sum l(i),\quad\text{where}$ (10)
@@
Figure 4 illustrates that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate such as $\alpha=10$, aligning with our expectations.
Refer to captionRefer to caption
Figure 4: Training curves for $\alpha=0.1$ and $\alpha=10$ with the Lipschitz continuous Huber loss function. As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$. This is to be expected from Proposition 5.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
@@
In section 3.1 of Ma et al. (2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near local minima, concluding that subquadratic growth explains the edge of stability phenomenon. However, their analysis is fundamentally flawed for these ReLU networks because second-order finite differences break down at their generically non-differentiable local minima (Laurent and Brecht, 2018). The overparameterized Huber loss function (Equation 10) further challenges the general applicability of their claims: despite having a global minimum of 0 with quadratic growth nearby, it exhibits unstable convergence as guaranteed by Proposition 5.1 (Figure 4). This discrepancy arises because the Huber loss lacks the twice differentiability required by Definition 4.1 in Ma et al. (2022). In contrast, all $L-smooth$ functions, being twice differentiable by definition, will show subquadratic growth near their local minima during unstable convergence, aligning with their results. This example underscores the risks of extending results derived for smooth functions to non-smooth cases.
@@
The limitations of applying results like those in Cohen et al. (2021) to non-smooth settings highlight the need for new theoretical frameworks. In hindsight, it is not particularly surprising that Cohen et al. (2021).’s results do not hold even for the differentiable case, as they derive their key result using the assumption of $L-smoothness$ while acknowledging in their discussion that this assumption is not ideal for neural networks. We propose using the convex non-differentiable single-layer neural network from Kumar (2023) as a toy problem for theory development. This approach offers advantages in terms of analytical tractability and relevance to ReLU networks, potentially leading to insights about optimization around non-differentiable points and informing more robust deep learning optimization techniques.